OpenStudy (anonymous):

Help me plz. Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (0, 0) m = 2

OpenStudy (senkai):

Well, y = mx + b. M is the slope, and it says m = 2. So, y = 2x + b. b is the y intercept, in (0,0). The last number is always the y-intercept. So, your answer would be y = 2x + 0, or y = 2x.

OpenStudy (anonymous):

Eq of line in slope point form is given as \[y-y_1=m(x-x_1) \] here x1=0, y1=0, m=2 (y-0)= 2(x-0) y=2x is the required eq of the line